# Astrophysical signals from the neighboring brane

### v2

@@ 5.3 Ultra-high-energy cosmic rays
Ultra-high-energy cosmic rays (UHECR) of the energy $10^{11}$eV [ParticleDataGroup:2024cfk] are one of the most intriguing mysteries of modern astrophysics. So far, only several protons of such energy were definitely registered. Despite numerous attempts to explain this phenomenon, its origin remains largely unknown. The discussion on the origin of the high-energy photons of PeV energy can be found in [2011APh....35...28B].
@@
This value is consistent with the lower boundary limits (4). The small number of such events is due to the extremely small cross-section, as previously discussed.
A similar mechanism used to explain high-energy photons encounters difficulties. The highest observed photon energy is about $10^{6}$ GeV, and the mass of a heavy electron must be of this order of the magnitude. However, this value is below the limit marked in (4), so some efforts should be applied to relate this effect to the dark matter.
@@ 6 Conclusion
Two-brane extra dimensions play an important role in theoretical research starting from the Randall-Sundrum model. In this research, we study the distribution of fields using the two-brane model [Popov:2024nax]. This model has revealed its ability to explain the electroweak hierarchy problem, the restoration of the Starobinsky inflationary model [Starobinsky:1980te], and the smallness of the cosmological constant [Bronnikov:2023lej].
Here we focus on the fields distribution over the extra dimensions. The extra-dimensional regions where fermions are localized are referred to as (thick) branes. It is shown that fermions split into two independent effective fields, localized on adjacent branes. This bifurcation process accompanies the evolution of the Universe from energy levels above the inflationary scale to the present-day state. The conditions under which the fermions and the Higgs field are independent at neighboring branes are attentively discussed in the subsection 4.2. At the same time,the gauge fields are uniformly distributed in the extra dimensions, preserving the charge universality in 4D physics.
The emergence of complex structures requires a hierarchy of energy scales and associated fine-tuning [2007unmu.book..231D]. In a brane-world model, the delicate process of parameter selection for this fine-tuning can be accomplished on a single brane, the one inhabited by the observers (brane-1). Consequently, a secondary brane (brane-2) would lack this fine-tuning, precluding the formation of stellar-like structures. This means that the masses of the particles on the other brane should have ’’natural’’ values, i.e. be of the order of the energy scale at which our Universe was formed, i.e. extremely massive. The Higgs mechanism of mass creation on both branes is discussed in subsection 5.2.
The stable matter located on the brane-2 most plausibly consists of the charged particles such as 3 generations of leptons and the protons as well as neutral particles like the neutrino and the atoms of hydrogen. Neutral particles are obvious candidates for the role of dark matter. However, the heavy-charged leptons are also promising candidates. The extremely high mass of such leptons on brane-2 is the key factor for the significant reduction of the Compton scattering cross section. This allows the heavy leptons to be considered as a component of dark matter.
It is shown that the fermions located on the different branes can exchange via photons and gravitons and thus be a source of ultra-high energy particles.
The Sandwich model of extra dimensions discussed here seems to be a promising framework that deserves further development. It could also serve as a basis for explaining phenomena such as the positron anomaly [Belotsky:2018vyt] and the smallness of the neutrino mass.
@@ Acknowledgements
The author is grateful to K. Belotsky and R. Konoplich for fruitful discussions. The work was partly funded by the Ministry of Science and Higher Education of the Russian Federation, Project "New Phenomena in Particle Physics and the Early Universe" FSWU-2023-0073. The research presented in Sect.3,4 was also partially carried out in accordance with the Kazan Federal University Strategic Academic Leadership Program.
[…]
[title=References, heading=bibintoc]

### v3

@@ 5.2 Ultra-high-energy cosmic rays
Ultra-high-energy cosmic rays (UHECR) of the energy $10^{11}$GeV [56] are one of the most intriguing mysteries of modern astrophysics. So far, only several protons of such energy were definitely registered. Despite numerous attempts to explain this phenomenon, its origin remains largely unknown. The discussion on the origin of the high-energy photons of PeV energy can be found in [16].
@@
This value is consistent with the lower boundary limits (47). Given the extremely small cross-section, such events occur rarely, as outlined earlier. For a rigorous analysis of superheavy charged particles in this framework, see [8]
A similar mechanism used to explain high-energy photons encounters difficulties. The highest observed photon energy is about $10^{6}$ GeV, and the mass of a heavy electron must be of this order of the magnitude. However, this value lies below the threshold defined in (47), so some efforts should be applied to connect this effect to the particles located in the uninhabited brane.
@@ 6 Conclusion
Two-brane extra dimensions play an important role in theoretical research starting from the Randall-Sundrum model.
In this research, we study the distribution of fields using the two-brane model developed in [61]. This model has revealed its ability to explain the electroweak hierarchy problem, the restoration of the Starobinsky inflationary model [70], and the smallness of the cosmological constant [24].
Here we focus on the fields distribution over the extra dimensions. The extra-dimensional regions where fermions are localized are referred to as (thick) branes. It is shown that fermions split into two independent effective fields, localized on adjacent branes. This bifurcation process accompanies the evolution of the Universe from energy levels above the inflationary scale to the present-day state. Therefore, the known Standard model particles are presented in both branes. The conditions under which the fermions and the Higgs field are independent at neighboring branes are attentively discussed in the subsection 4.2. At the same time, the gauge fields are uniformly distributed in the extra dimensions, preserving the charge universality in 4D physics.
The emergence of complex structures requires a hierarchy of energy scales and associated fine-tuning [35]. In a brane-world model, the delicate process of parameter selection for the fine-tuning can be accomplished on a single brane, the one inhabited by the observers (brane-1). This means that the masses of the particles, partners of Standard Model particles in the other brane should have ’’natural’’ values, i.e. be of the order of the energy scale at which our Universe was formed, i.e. extremely massive. The Higgs mechanism of mass creation on both branes is discussed in subsection 4.6. In addition, a secondary brane (brane-2) are lack the fine-tuning, precluding the formation of stellar-like structures.
The stable matter located on the brane-2 consists of the charged particles such as 3 generations of leptons and the protons as well as neutral particles like the neutrinos and the atoms of hydrogen. Neutral particles are obvious candidates for the role of dark matter. The extremely high mass of heavy charged leptons on brane-2 is the key factor for the significant reduction of the Compton scattering cross section. This allows the heavy charged leptons to be considered as a component of dark matter, albeit with a likely small contribution due to stringent observational constraints.
It is also shown that the fermions located on the different branes can exchange via photons and gravitons and thus be a source of ultra-high energy particles.
The model of extra dimensions discussed here seems to be a promising framework that deserves further development. It could also serve as a basis for explaining phenomena such as the positron anomaly [15] and the smallness of the neutrino mass.
@@ Acknowledgements
The author is grateful to K. Belotsky and R. Konoplich for fruitful discussions. The work was partly funded by the Ministry of Science and Higher Education of the Russian Federation, Project "Studying physical phenomena in the micro- and macro-world to develop future technologies" FSWU-2026-0010. The research presented in Sect.3,4 was also partially carried out in accordance with the Kazan Federal University Strategic Academic Leadership Program.
Conflicts of interests
The author confirm that there are no any potential conflict(s) of interest like employment, consulting fees, research contracts, stock ownership, patent licenses, honoraria, advisory affiliations etc.